OpenStudy (anonymous):

Benito made a wooden box to grow herbs on his patio. The box is 16 inches wide, 12 inches tall, and 28 inches long. He plans to fill the box exactly two-thirds full with potting soil. What volume of soil does Benito need?

OpenStudy (butterflydreamer):

Firstly, find the total volume of the box which can be filled with soil:) To do this Volume (V) = 16 * 12 * 28 = ? Once we know this, since the question tells you that the person wants to fill up only 2/3 of the box with soil. You would do 2/3 * volume of box = ? Basically, find the volume of the box (as this is how much soil can fill the ENTIRE box) and since we only want to fill up 2/3 of the box with soil, you would multiply the volume of the box by 2/3 and you'll get your answer :)

OpenStudy (butterflydreamer):

@Vocaroo